OpenStudy (anonymous):

find a formula for the inverse f(x)=x+9

OpenStudy (anonymous):

switch x and y, then solve for y

OpenStudy (anonymous):

swap f(x) and x. then solve for f(x).

OpenStudy (anonymous):

x=y+9 y=x-9

OpenStudy (anonymous):

x=y+9 y=x-9 inverse f(x)=x-9